OpenStudy (darkigloo):

How do I find the domain of √(x^4 - 16x^2) ?

OpenStudy (ranga):

We cannot take the square root of a negative number. Therefore, x^4 - 16x^2 >= 0 x^2(x^2 - 16) >= 0 x^2 is always >= 0 Therefore x^2 - 16 must be >= 0 solve this inequality and that will give you the domain.

OpenStudy (darkigloo):

Ah I see. so x>=4 and x=<-4 and x=0?

OpenStudy (ranga):

x <= -4 OR x >= +4. (not x = 0). In interval notation, the domain is: (-infinity, -4] U [4, infinity) where U is the union symbol.

OpenStudy (darkigloo):

can you please show me step by step how to get the answer?

OpenStudy (ranga):

The first reply has the step-by-step solution.

OpenStudy (darkigloo):

oh alright. thank you. :)

OpenStudy (ranga):

you are welcome.